OpenStudy (anonymous):

I NEED HELP! PLEASE SOMEONE HELP ME OUT! tomorrow's my last day to submit my assignment. A man is currently 4 times as old as his son. In 4 years from now he will be three times as old as his son will be then. How old is his son now?

OpenStudy (anonymous):

Let the age of man be m, and the age of son be s. Currently 4s=m In 4 years 3(s+3)=(m+3) Subtract the second equation from the first to get s-9=-3 so s=6 m=6*4=24 Son is 9 years old now, dad is 27.

OpenStudy (akshay_budhkar):

o.O you dont satisfy the original equation andreas

OpenStudy (anonymous):

Oops this is wrong :) 3(s+3)=(m+3) should be 3(s+4)=(m+4) So s-12=-4 s=8 m=32